If a hydrocarbon sample is burned in air to produce 0.22 grams of CO2, what is the mass of carbon in the sample?
1
Expert's answer
2019-06-07T03:16:08-0400

The mass of carbon in the sample is the same as the mass of carbon in the carbon dioxide produced.

"n(CO_2)=\frac{m(CO_2)}{M(CO_2)}=\frac{m(C \ in \ CO_2)}{M(C)}"

"m(C \ in \ CO_2) = m(CO_2)\frac{M(C)}{M(CO_2)}=0.22g \cdot \frac{12\frac{g}{mol}}{44\frac{g}{mol}}=0.06g"

Thus the mass of carbon in the sample is 0.06g